The demand curve function of a monopolist is given by P= 100 - 2Q. If the marginal cost is constant and is equal to 20, what is the amount of profit made?
1
Expert's answer
2020-12-15T07:22:16-0500

Solution:

"P=100-2Q"

Derive MR:

"R=(100)\times (Q)- (2Q)(Q) = 100Q-2Q^{2}"


"MR=\frac{dR}{dQ} = 100-4Q"


"MC =20"

"MR=MC"


"100-4Q=20"


"4Q=100-20"

"4Q=80"


"Q=\frac{80}{4} = 20"

"Q=20"

Substitute to get the price:

"P=100-(2)\times (20) =100-40=60"

"P=60"


Derive Total Revenue (TR) and Total Cost (TC):

"TR=(60\; per\; unit)\times (20 \;units\; per\; unit) = 1,200"


"TC=(10\; per\; unit)\times (20 \;units\; per\; unit) = 200"


"Profit=TR-TC=1,200-200 = 1,000"


"Profit \;made = 1,000"



Profit made = 1,000